OpenStudy (anonymous):

Can someone please help me with these two quadratic equations? I kind of understand how to do it but I can't seem to get the right answers. 1. Solve 2x^2 - 8x = -7 2. If a baseball player hits a baseball from 4 feet off the ground with an initial velocity of 64 feet per second, how long will it take the baseball to hit the ground? Use the equation h = -16t2 + 64t + 4.

OpenStudy (phi):

2x^2 - 8x = -7 put into standard form by add + 7 to both sides. you will get 2 x^2 -8 x + 7=0 now use the quadratic formula. see http://www.khanacademy.org/math/trigonometry/polynomial_and_rational/quad_formula_tutorial/v/using-the-quadratic-formula

OpenStudy (anonymous):

I know the formula. I always have trouble plugging the numbers into the formula.

OpenStudy (phi):

for the problem 2 x^2 -8 x + 7=0 what is a, b and c ?

OpenStudy (anonymous):

A=2 B=-8 and C=7 is that correct?

OpenStudy (phi):

so far, so good now write down the quadratic formula with the letters then underneath it, write it again, but replace the letters with numbers don't do any arithmetic. just write it down.

OpenStudy (anonymous):

\[\frac{ -b+-\sqrt{b^2-4AC} }{ 2a }\]

OpenStudy (phi):

looks good. now replace the letters with the numbers A=2 B=-8 and C=7

OpenStudy (anonymous):

\[\frac{ -8+-\sqrt{-8^2-4(2)(7)} }{ 2(2) }\]

OpenStudy (phi):

to do this part correctly, we have to be careful of the minus signs the formula wants -b b= -8 put b in parens, like this: (-8) now use that to replace b in the formula -b becomes -(-8) notice that is not what you have. -(-8) is the same as +8 or just 8

OpenStudy (phi):

also, inside the square root, write b^2 as (-8)^2 (which means -8 *-8)

OpenStudy (anonymous):

Is this correct? \[\frac{ -(-8)+-\sqrt{-8^2-4(2)(7)} }{ 2(2) }\]

OpenStudy (phi):

almost, inside the square root, write it (-8)^2

OpenStudy (anonymous):

Is that right now lol? \[\frac{ -(-8)+-\sqrt{(-8)^2-4(2)(7)} }{ 2(2) }\]

OpenStudy (phi):

perfect. now let's concentrate on the stuff inside the square root can you simplify the (-8)^2 - 4(2)(7) ?

OpenStudy (anonymous):

I'll be right back.. please don't leave, it'll only take me 20 minutes

OpenStudy (phi):

message me when you get back

OpenStudy (anonymous):

okay I am back, thank you for waiting. I simplified whats inside the square root and got \[2\sqrt{2}\] Is that correct?

OpenStudy (phi):

yes. almost done simplify the -(-8) out front to + 8 you now have \[ \frac{8 ± 2 \sqrt{2}}{2\cdot 2} \]

OpenStudy (phi):

yes, btw in the equation editor if you click on the top row, it brings up a list of operators, one of which is ± but to continue. you can divide top and bottom by 2 to get one form of the answer

OpenStudy (anonymous):

can you show me how the equation should look when you divide the top and bottom. I don't think i did it right. I got \[\frac{ 1 }{ 4 }(4+\sqrt{2}\]

OpenStudy (anonymous):

my answer options are A. \[-2\pm \sqrt{2} \] b. \[-2\pm2\sqrt{2}\] c. \[\frac{ 2\pm \sqrt{2} }{ 2 }\] d.\[2\pm \frac{ \sqrt{2} }{ 2 }\]

OpenStudy (phi):

remember when adding or subtracting fractions you need a common denominator? for example \[ \frac{3}{4} - \frac{1}{4} = \frac{3 -1}{4} \] notice we can "go backward" and write \[ \frac{3 -1}{4}=\frac{3}{4} - \frac{1}{4} \] use that same idea with \[ \frac{8 ± 2 \sqrt{2}}{2\cdot 2} = \frac{8}{2\cdot 2}± \frac{2\sqrt{2}}{2\cdot 2}\]

OpenStudy (phi):

now you have two separate divide problems 8/4 and the other term

OpenStudy (anonymous):

I got \[\frac{ 8 }{ 2*2 }=2 & \frac{ 2\sqrt{2} }{ 2*2 }=\frac{ \sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ 8 }{ 2*2 }=2 \]

OpenStudy (phi):

looks correct.

OpenStudy (anonymous):

and \[\frac{ 2\sqrt{2} }{ 2*2 }=\frac{ \sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

so the answer would be \[2\pm \frac{ \sqrt{2} }{ 2 }\] and be D.

OpenStudy (phi):

yes

OpenStudy (anonymous):

Also, even though I HATE not doing my own work can you pretty please just give me the answer to --> If a baseball player hits a baseball from 4 feet off the ground with an initial velocity of 64 feet per second, how long will it take the baseball to hit the ground? Use the equation h = -16t2 + 64t + 4 a. \[2\pm \frac{ \sqrt{17} }{ 2 }\] b.\[\frac{ 2\pm \sqrt{17} }{ 2 }\] c. \[2\pm4\sqrt{17}\] d.\[\frac{ 16\pm \sqrt{17} }{ 2 }\]

OpenStudy (phi):

h = -16t^2 + 64t + 4 the height h = 0 when on the ground so you must solve -16t^2 + 64t + 4 =0 the first thing you can do is divide both sides (all terms by 4) to simplify it -4 t^2 + 16 t + 1 = 0 now use the quadratic formula.

OpenStudy (anonymous):

actually that is all I need. Word problems aren't my strong. But after you got it looking right and all the help you already gave me its no problem. THANK YOU! You have no idea how much you helped.